# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

### v1

@@ Abstract
In this study, we proposed a design methodology for a piezoelectric energy-harvesting device optimized for maximal power generation at a designated frequency using topology optimization. The proposed methodology emphasizes the design of a unimorph-type piezoelectric energy harvester, wherein a piezoelectric film is affixed to a singular side of a silicon cantilever beam. Both the substrate and the piezoelectric film components underwent concurrent optimization. Constraints were imposed to ensure that the resultant design is amenable to microfabrication, with specific emphasis on the etchability of piezoelectric energy harvesters. Several numerical examples were provided to validate the efficacy of the proposed method. The results showed that the proposed method derives both the substrate and piezoelectric designs that maximize the electromechanical coupling coefficient and allows the eigenfrequency of the device and minimum output voltage to be set to the desired values. Furthermore, the proposed method can provide solutions that satisfy the cross-sectional shape, substrate-depend, and minimum output voltage constraints. The solutions obtained by the proposed method are manufacturable in the field of microfabrication.
@@ 1 Introduction
Piezoelectric energy harvesters convert mechanical energy into electrical energy through the piezoelectric effect. These devices have gained significant attention in recent years owing to their potential to scavenge energy from various sources such as vibrations Jeon et al. (2005); Roundy et al. (2003) and fluid flow Yun and Kim (2016). The design of a piezoelectric energy harvester plays a critical role in determining its efficiency and performance Erturk and Inman (2011). Various piezoelectric energy harvester designs, including cantilever beams, multilayered structures, and cylinder structures, have been proposed and studied Cook-Chennault et al. (2008); Beeby et al. (2006). The unimorph cantilevered piezoelectric energy harvester has several advantages over the other piezoelectric energy harvester designs. The simple and compact design of this harvester makes it suitable for applications where size and weight are critical factors, such as in portable or wearable devices. Furthermore, the unimorph cantilevered design allows easy integration with other systems such as sensors or wireless transmitters Saxena et al. (2021). However, the performance of a unimorph cantilevered piezoelectric energy harvester is limited by the amount of mechanical energy harvested from external vibrations. To overcome these limitations, researchers have explored the use of advanced materials and optimization techniques to improve the performance of unimorph cantilevered piezoelectric energy harvesters. Some studies have investigated piezoelectric materials with improved electromechanical coupling factors to enhance the piezoelectric effect and improve energy conversion efficiency Yoshimura et al. (2013); Murakami et al. (2013). Other researchers have focused on designing the geometry of the cantilever beam and piezoelectric element for maximum energy output. For instance, rectangular, triangular, and trapezoidal shapes have been proposed for unimorph piezoelectric vibration energy harvesters Glynne-Jones et al. (2001). However, as these structures have simple geometries, significant improvements can be achieved using topological optimization techniques Bendsøe and Kikuchi (1988).
@@
Sato et al. Sato et al. (2017), Yamada et al. Yamada and Noguchi (2022), and Tajima et al. Tajima and Yamada (2023) developed advanced methodologies that leverage fictitious physical models to assess manufacturability. Sato et al. Sato et al. (2017) developed a method for evaluating manufacturability in molding processes using fictitious physical models to ensure that the geometrical features of mold parts are amenable to manufacturing. Yamada et al. Yamada and Noguchi (2022) focused on additive manufacturing and proposed a topology optimization method that addressed closed cavity exclusion constraints and enhanced the design process of additive manufacturing. Tajima et al. Tajima and Yamada (2023) introduced a coupled fictitious physics model aimed at improving the convergence in topology optimization by integrating the objectives of both the fictitious physical and mechanical models, particularly for additive manufacturing with geometric constraints. These methodologies collectively underscore the importance of incorporating manufacturability considerations into topology optimization, which is essential for practical applications in various manufacturing processes. The level set-based topology optimization championed by Yamada et al. Yamada et al. (2010) affords precision in modulating the geometrical intricacy of the resultant optimal configurations. Building on this foundation, Noda et al. Noda et al. (2022) augmented this level set strategy to cater to the nuances of multi-material topology optimization. This innovation ensures a meticulous orchestration of the complexity intrinsic to each deployed material, culminating in a multi-material assembly that is conducive to straightforward manufacturing. However, the previously underscored methodologies are predominantly conceptualized for traditional, macroscopic machining techniques such as milling. Therefore, their direct application to piezoelectric energy harvesters, typically realized through microfabrication techniques such as etching Murakami et al. (2013); Aramaki et al. (2019), may not be feasible.
@@
In this study, we introduced a methodology for the optimal design of piezoelectric energy harvesters by employing topology optimization anchored in the level set method. This approach emphasizes manufacturability in the field of microfabrication, with etching being the predominant process under consideration. The design criteria for piezoelectric energy harvesters were delineated based on the operational frequency and a threshold for the excitation voltage output. Furthermore, the proposed method uniquely facilitates the concurrent optimization of three-dimensional structures within both the substrate and piezoelectric material domains. The resulting designs not only satisfy the stipulated specifications but are also readily adaptable to microfabrication procedures.
@@
The remainder of this paper is organized as follows: In Section 2, the architectural framework of the designated device is described. After this initial exposition, in this section we further elaborate on the optimal design approach, articulating the pivotal design criteria, including the eigenfrequency specifications and constraints regarding the minimum output voltage. In Section 3, we explain the formulation of constraints pertinent to manufacturability. Specifically, we introduced two constraints to enhance manufacturability: ensuring a consistent cross-sectional shape across each domain and the stipulation that no piezoelectric material is positioned without an accompanying substrate. In Section 4, implementation of the numerical analysis of the proposed method is described. In Section 5, numerical illustrations by employing a benchmark model to test the efficacy of the introduced method are detailed. Finally, in Section 6, the conclusions of this study are summarized. Additionally, the methodologies tailored for computing the output voltage are detailed in A, and the sensitivity analyses are discussed in B.
@@ 2.1 Unimorph cantilevered piezoelectric energy harvester
This study focused predominantly on the unimorph cantilevered energy harvester. An unimorph cantilevered piezoelectric energy harvester comprises three main components: substrate, piezoelectric material, and weight. These elements are depicted in Figure 1, which shows the theoretical structure of an unimorph cantilevered piezoelectric vibration energy harvester. The unique advantage of the unimorph design is its capability to produce a superior electrical output for a given mechanical input compared with a conventional cantilevered piezoelectric energy harvester. This superiority is attributed to the mechanical amplification effect induced by the nonpiezoelectric layer, which result in an enhanced deformation of the piezoelectric layer for a specific input vibration.
@@ 2.2 Formulation of piezoelectric phenomena
The governing equations for a piezoelectric energy harvester are articulated in the following set of simultaneous piezoelectric equations, as elaborated in Erturk and Inman (2011):
$\begin{split}\bm{\sigma}&=\bm{C}^{E}\bm{s}(\bm{u})+\bm{e}^{T}\bm{\nabla}\varphi\\ \bm{d}&=\bm{e}\bm{s}(\bm{u})+\bm{\varepsilon}^{S}\bm{\nabla}\varphi,\end{split}$ (1)
$\bm{s}(\bm{u})=\frac{1}{2}(\bm{\nabla}\bm{u}+(\bm{\nabla}\bm{u})^{T}).$ (2)
@@
Here, $\bm{\sigma}$ is the stress tensor; $\bm{d}$ is the electric displacement tensor; $\bm{u}$ is the displacement vector; $\bm{s}(\bm{u})$ represents the strain tensor; $\bm{C}^{E}$ is the elasticity tensor at a constant electric field; and we assume that $\bm{C}^{E}$ represents an isotropic linear elastic material; $\bm{e}$ represents the piezoelectric stress constant tensor; $\varphi$ is the electric potential; $\bm{\varepsilon}^{S}$ is the dielectric constant tensor at a constant strain; and ${}^{T}$ denotes the transpose of the tensor.
@@
In this study, we primarily focused on a piezoelectric energy harvester under harmonic excitation at a frequency $\omega$. Within this context, the displacement vector $\bm{u}$ and the electric potential $\varphi$ are designated as state variables, whereas $\bm{v}$ and $v_{\varphi}$ serve as test functions. Consequently, Eq. (1) is transcribed into its weak form as follows:
$\begin{split}\int_{\Omega}\bm{s}(\bm{v})^{T}\bm{C}^{E}\bm{s}(\bm{u})\ \text{d}\Omega-\omega^{2}\rho\int_{\Omega}\bm{v}^{T}\bm{u}\ \text{d}\Omega+\int_{\Omega}\bm{s}(\bm{v})^{T}\bm{e}^{T}\bm{\nabla}\varphi\ \text{d}\Omega=\int_{\Omega}\bm{v}^{T}\hat{\bm{f}}\ \text{d}\Omega\\ \int_{\Omega}(\bm{\nabla}v_{\varphi})^{T}\bm{e}\bm{s}(\bm{u})\ \text{d}\Omega-\int_{\Omega}(\bm{\nabla}v_{\varphi})^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi\ \text{d}\Omega=\int_{\Omega}v_{\varphi}\hat{\bm{q}}\ \text{d}\Omega.\end{split}$ (3)
@@
Here, $\rho$ refers to the material’s density, $\hat{\bm{f}}$ is the external forcing function, and $\hat{\bm{q}}$ represents the externally provided electric charge. In this study, we assumed steady-state vibration.
The electromechanical coupling coefficient was employed as a metric to assess the performance of the piezoelectric energy harvester. The coefficient is calculated using the following equation, which considers the $i$-th mode eigenfrequency $\omega_{oci}$ under electrode open-circuit conditions and the $i$-th mode eigenfrequency $\omega_{sci}$ under electrode short-circuit conditions Hagood and von Flotow (1991):
@@
Under open-circuit conditions (where $\hat{\bm{q}}=0$), the piezoelectric simultaneous equations are expressed as follows:
$\begin{split}\int_{\Omega}\bm{s}(\bm{v}_{oci})^{T}\bm{C}^{E}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\omega_{oci}^{2}\rho\int_{\Omega}\bm{v}_{oci}^{T}\bm{u}_{oci}\ \text{d}\Omega+\int_{\Omega}\bm{s}(\bm{v}_{oci})^{T}\bm{e}^{T}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega&=0\\ \int_{\Omega}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{e}\bm{s}(\bm{u}_{oci})\ \text{d}\Omega-\int_{\Omega}\bm{\nabla}v_{\varphi_{oci}}^{T}\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}\ \text{d}\Omega&=0,\end{split}$ (5)
where $\bm{u}_{oci}$ is the eigenvector corresponding to the eigenvalue $\omega_{oci}$, $\varphi_{oci}$ denotes electric potential under open-circuit condition, and $\bm{v}_{oci}$ is the test function. Conversely, under short-circuit conditions (where $\bm{\nabla}\varphi=0$), the piezoelectric simultaneous equations become solely the equation for harmonic oscillation, unaffected by the influence of the piezoelectric material.
$\int_{\Omega}\bm{s}(\bm{v}_{sci})^{T}\bm{C}^{E}\bm{s}(\bm{u}_{sci})\ \text{d}\Omega-\omega_{sci}^{2}\rho\int_{\Omega}\bm{v}_{sci}^{T}\bm{u}_{sci}\ \text{d}\Omega=0,$ (6)
where $\bm{u}_{sci}$ is the eigenvector corresponding to the eigenvalue $\omega_{sci}$ and $\bm{v}_{sci}$ is the test function.
In the context of piezoelectric energy harvester design, the eigenfrequency of the device is established as a crucial design parameter. Consequently, in this study, we considered strategies for concurrently aligning the eigenfrequency with the desired value. Adopting the objective function proposed by Yamada et al. Yamada et al. (2012), we normalized the eigenfrequency specification problem with respect to the $i$-th mode desired frequency $\bar{\omega}_{i}$ and formulated the objective function $F{\omega}$ as follows:
@@
where we considered $F_{\omega}$ as a problem specifying up to the $n$-th mode.
To transform the maximization problem into a minimization problem, we formulated the objective function $F_{k}$ using the electromechanical coupling coefficient $k_{i}$ as follows:

### v2

@@ Abstract
In this study, we propose a design methodology for a piezoelectric energy-harvesting device optimized for maximal power generation at a designated frequency using topology optimization. The proposed methodology is adapted to the design of a unimorph-type piezoelectric energy harvester, wherein a piezoelectric film is affixed to a singular side of a silicon cantilever beam. Both the substrate and the piezoelectric film components undergo concurrent optimization. Constraints are imposed to ensure that the resultant design is amenable to microfabrication, with specific emphasis on the etchability of piezoelectric energy harvesters. Several numerical examples are provided to validate the efficacy of the proposed method. The results show that the proposed method yields optimized substrate and piezoelectric designs with an enhanced electromechanical coupling coefficient, while allowing the eigenfrequency of the device and the minimum output voltage to be set to the desired values. Furthermore, the proposed method can provide solutions that satisfy the cross-sectional shape, substrate-dependent, and minimum output voltage constraints. The solutions obtained by the proposed method are manufacturable in the field of microfabrication.
@@ 1 Introduction
Piezoelectric energy harvesters convert mechanical energy into electrical energy through the piezoelectric effect. These devices have gained significant attention in recent years owing to their potential to scavenge energy from various sources such as vibrations [1, 2] and fluid flow [3]. The design of a piezoelectric energy harvester plays a critical role in determining its efficiency and performance [4]. Various piezoelectric energy harvester designs, including cantilever beams, multilayered structures, and cylindrical structures, have been proposed and studied [5, 6]. The unimorph cantilevered piezoelectric energy harvester has several advantages over other piezoelectric energy harvester designs. A unimorph cantilevered piezoelectric energy harvester consists of a cantilever beam with one end fixed and the other end free to move. The beam is typically composed of two primary domains: a substrate domain and a piezoelectric material domain. Additionally, it includes electrode domains for electrical energy extraction. When the device is subjected to external vibrations or mechanical loads, the beam bends, causing mechanical strain in the piezoelectric domain. This strain generates electrical energy through the piezoelectric effect. The simple and compact design of this harvester makes it suitable for applications where size and weight are critical factors, such as in portable or wearable devices. Furthermore, the unimorph cantilevered design allows easy integration with other systems, such as sensors or wireless transmitters [7]. However, the performance of a unimorph cantilevered piezoelectric energy harvester is limited by the amount of mechanical energy harvested from external vibrations. To overcome these limitations, researchers have explored the use of advanced materials and optimization techniques to improve their performance. Some studies have investigated piezoelectric materials with improved electromechanical coupling factors to enhance the piezoelectric effect and improve the efficiency of energy conversion [8, 9]. Other researchers have focused on designing the geometry of the cantilever beam and piezoelectric element for maximum energy output. For instance, rectangular, triangular, and trapezoidal shapes have been proposed for unimorph piezoelectric vibration energy harvesters [10]. However, as these structures have simple geometries, significant improvements can be achieved using topological optimization techniques [11].
@@
However, these methodologies often yield designs that face inherent manufacturing challenges. Consequently, various strategies have been proposed to realize manufacturable designs through topology optimization. Vatanabe et al. [26], Michailidis et al. [27], and Sato et al. [28] developed advanced methodologies to assess manufacturability. Vatanabe et al. [26] proposed a unified projection-based approach to incorporate manufacturing constraints into topology optimization, enabling the generation of manufacturable designs with minimum member size, minimum hole size, symmetry, and other geometric constraints. This method ensures that the optimized structures are not only optimal in terms of performance but also feasible for manufacturing processes like casting, milling, turning, forging, and rolling. Michailidis et al. [27] focused on multi-phase shape and topology optimization via a level-set method, incorporating manufacturing constraints to ensure feasible designs for practical applications. Sato et al. [28] developed a method for evaluating manufacturability in molding processes using fictitious physical models to ensure that the geometrical features of molded parts are amenable to manufacturing. Building upon the advancements in manufacturability assessments, various studies have proposed constraints related to manufacturability specifically in the context of additive manufacturing. Liu et al. [29] reviewed the current trends and future directions in topology optimization for additive manufacturing, highlighting the need for support structure optimization and addressing issues related to overhangs and material anisotropy. Yamada et al. [30] focused on additive manufacturing and proposed a topology optimization method that addressed closed cavity exclusion constraints and enhanced the design process of additive manufacturing. Tajima et al. [31] introduced a coupled fictitious physics model aimed at improving the convergence in topology optimization by integrating the objectives of both the fictitious physical and mechanical models, particularly for additive manufacturing with geometric constraints. These methodologies collectively underscore the importance of incorporating manufacturability considerations into topology optimization, which is essential for practical applications in various manufacturing processes. The level set-based topology optimization championed by Yamada et al. [32] affords precision in modulating the geometrical intricacy of the resultant optimal configurations. Moreover, earlier studies developed various methodologies for multi-material topology optimization, such as those by Yulin and Xiaoming [33] and Wang and Wang [34], which laid the groundwork for subsequent advancements in this field. Building on this foundation, Noda et al. [35] augmented the level set method to address the specific challenges of multi-material topology optimization. This innovation ensures a meticulous orchestration of the complexity intrinsic to each deployed material, culminating in a multi-material assembly that is conducive to straightforward manufacturing. However, the previously underscored methodologies are predominantly conceptualized for traditional, macroscopic machining techniques such as milling. Therefore, their direct application to piezoelectric energy harvesters, typically realized through microfabrication techniques such as etching [9, 36], may not be feasible.
@@
In this study, we introduce a methodology for the optimal design of piezoelectric energy harvesters by employing topology optimization based on the level set method. This approach emphasizes manufacturability in the field of microfabrication, with etching being the predominant process under consideration. The design criteria for piezoelectric energy harvesters are delineated based on the operational frequency and a threshold for the excitation voltage output. Furthermore, the proposed method uniquely facilitates the concurrent optimization of three-dimensional structures within both the substrate and piezoelectric material domains. The resulting designs not only satisfy the stipulated specifications but are also readily adaptable to microfabrication procedures.
@@
The remainder of this paper is organized as follows: In Section 2, the architectural framework of the designated device is described. After this initial exposition, in that section we further elaborate on the optimal design approach, articulating the pivotal design criteria, including the eigenfrequency specifications and constraints regarding the minimum output voltage. In Section 3, we explain the formulation of constraints pertinent to manufacturability. Specifically, we introduce two constraints to enhance manufacturability: ensuring a consistent cross-sectional shape across each domain and stipulation that no piezoelectric material is positioned without an accompanying substrate. In Section 4, the implementation of the proposed method is described. In Section 5, numerical illustrations employing a benchmark model to test the efficacy of the introduced method are detailed. Finally, in Section 6, the conclusions of this study are summarized. Additionally, the methodologies tailored for computing the output voltage are detailed in A, and the sensitivity analyses are discussed in B.
@@ 2.1 Unimorph cantilevered piezoelectric energy harvester
This study predominantly focuses on the unimorph cantilevered energy harvester. A unimorph cantilevered piezoelectric energy harvester comprises three main components: substrate, piezoelectric material, and weight. These elements are depicted in Figure 1, which shows the theoretical structure of a unimorph cantilevered piezoelectric vibration energy harvester. A unimorph cantilevered piezoelectric energy harvester consists of a cantilever beam with one end fixed and the other end free to move. The beam has a structure typically composed of three primary domains: a substrate domain, a piezoelectric material domain, and a weight domain at the free end. Additionally, it includes electrode domains for electrical energy extraction. The substrate domain is nonpiezoelectric and serves as the base layer to which the piezoelectric material is bonded. In this study, the substrate is composed of silicon, and the piezoelectric material utilized is lead zirconate titanate (PZT). To ensure electrical insulation between the electrodes and the substrate domain, an insulating layer is formed on both the upper and lower surfaces of the substrate domain. When the device is subjected to external vibrations or mechanical loads, the weight causes the beam to bend more significantly, inducing mechanical strain in the piezoelectric domain. This strain generates electrical energy through the piezoelectric effect. The presence of the weight enhances the bending of the beam, resulting in greater deformation of the piezoelectric layer and thus a higher electrical output.
The piezoelectric effect is a phenomenon where certain materials, such as quartz and PZT, generate an electric charge in response to applied mechanical stress. In a piezoelectric material, when mechanical strain is induced, an electric potential is developed across the material due to the displacement of charge carriers. This property allows piezoelectric materials to be used in energy harvesting devices, converting mechanical energy from vibrations into electrical energy.
The unique advantage of the unimorph design is its capability to produce a superior electrical output for a given mechanical input compared to a conventional cantilevered piezoelectric energy harvester. This superiority is attributed to the mechanical amplification effect induced by the substrate domain, which results in an enhanced deformation of the piezoelectric layer for a given input vibration. From a design viewpoint, the unimorph-type harvester features a simpler structure with a single piezoelectric layer, allowing for easier fabrication and reduced manufacturing costs. The design flexibility is also increased as both the substrate and the piezoelectric layer can be independently etched, allowing for more complex geometries and optimized vibration characteristics.
@@ 2.2 Formulation of piezoelectric phenomena
The governing equations for a piezoelectric energy harvester are articulated in the following set of simultaneous piezoelectric equations, as elaborated in [4]:
$\begin{split}\sigma_{ij}&=C^{E}_{ijkl}s_{kl}+e_{kij}\varphi_{,k}\\ d_{i}&=e_{ikl}s_{kl}-{\varepsilon}^{S}_{ik}\varphi_{,k},\end{split}$ (1)
@@
Here, $\sigma$ is the stress tensor; $d$ is the electric displacement tensor; $s$ represents the strain tensor; ${C}^{E}$ is the elasticity tensor at a constant electric field, and we assume that $C^{E}$ represents an isotropic linear elastic material; $e$ represents the piezoelectric stress constant tensor; $\varphi$ is the electric potential; $\varphi_{,k}$ is a derivative with respect to the $k$ th variable; and ${\varepsilon}^{S}$ is the dielectric constant tensor at a constant strain.
The piezoelectric coefficient tensor $e$ plays a crucial role in these equations, as it describes the coupling between the mechanical and electrical states of the material. In the second equation, $e_{kij}$ represents the piezoelectric effect, where mechanical strain $s_{kl}$ induces an electric displacement $d_{i}$. Conversely, in the first equation, the application of an electric field $\varphi_{,k}$ results in mechanical stress $\sigma_{ij}$. This bidirectional coupling is the hallmark of piezoelectric materials, allowing them to convert mechanical energy into electrical energy and vice versa. Therefore, the piezoelectric effect creates a strong interaction between the electric field and the elastic field within the material, enabling it to function efficiently as an energy harvester, sensor, or actuator.
2.2.1 Matrix and vector formulation of piezoelectric phenomena
To facilitate numerical computations and implementation, it is often convenient to convert these tensor equations into matrix and vector forms. This involves representing the strain tensor $s$ in its vector form, commonly using Voigt notation, and expressing the piezoelectric stress constant tensor $e$ and other relevant tensors in matrix forms.
Using Voigt notation, the strain tensor $s$ can be represented as a column vector $\bm{s}(\bm{u})$, and the elasticity tensor ${C}^{E}$ and piezoelectric stress constant tensor $e$ are transformed accordingly. The electric field $E_{k}$ is related to the electric potential $\varphi$ by $E_{k}=-\varphi_{,k}$. Consequently, the piezoelectric equations in matrix-vector forms are given by:
$\begin{split}\bm{\sigma}&=\bm{C}^{E}\bm{s}(\bm{u})+\bm{e}^{T}\bm{\nabla}\varphi\\ \bm{d}&=\bm{e}\bm{s}(\bm{u})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi.\end{split}$ (2)
In these equations, $\bm{\sigma}$ represents the stress vector (6$\times$1); $\bm{d}$ is the electric displacement vector (3$\times$1); $\bm{s}(\bm{u})$ is the strain vector (6$\times$1); $\bm{C}^{E}$ is the elasticity matrix (6$\times$6); $\bm{e}$ is the piezoelectric coefficient matrix (3$\times$6); $\bm{\varepsilon}^{S}$ is the dielectric constant matrix (3$\times$3); and $\bm{\nabla}\varphi$ is the electric field vector (3$\times$1) derived from the electric potential $\varphi$.
@@ 2.2.1 Matrix and vector formulation of piezoelectric phenomena
In this study, we primarily focus on the steady-state vibration of a piezoelectric energy harvester under harmonic excitation at a frequency $\omega$. Within this context, the displacement vector $\bm{u}$ and the electric potential $\varphi$ are designated as state variables.
Assuming a time-harmonic form for the displacement and electric potential, such that $\bm{u}(t)=\bm{u}e^{i\omega t}$ and $\varphi(t)=\varphi e^{i\omega t}$, the time-dependent piezoelectric equations with external forcing can be expressed as follows:
$\begin{split}\bm{C}^{E}\bm{s}(\bm{u})+\bm{e}^{T}\bm{\nabla}\varphi&=\rho\frac{\partial^{2}\bm{u}}{\partial t^{2}}+\hat{\bm{f}}\\ \bm{e}\bm{s}(\bm{u})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi&=\hat{\bm{q}}.\end{split}$ (3)
@@
Here, $\rho$ denotes the material’s density, $\hat{\bm{f}}$ represents the external forcing function, and $\hat{\bm{q}}$ indicates the externally provided electric charge.
By substituting $\bm{u}(t)=\bm{u}e^{i\omega t}$ and $\varphi(t)=\varphi e^{i\omega t}$ into these equations, we obtain the following:
$\begin{split}\bm{C}^{E}\bm{s}(\bm{u})-\omega^{2}\rho\bm{u}+\bm{e}^{T}\bm{\nabla}\varphi=\hat{\bm{f}}\\ \bm{e}\bm{s}(\bm{u})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi=\hat{\bm{q}}.\end{split}$ (4)
Note that the term $-\omega^{2}\rho\bm{u}$ arises from the second derivative of the harmonic displacement $\bm{u}e^{i\omega t}$ with respect to time, representing the inertia force.
To determine the eigenfrequencies of the system, we consider the free vibration case with no external forces or charges. This leads to the eigenvalue problem of the piezoelectric field under open-circuit and short-circuit conditions. In the open-circuit condition, the top and bottom electrodes are not connected, implying the electrical resistance between the electrodes is assumed to be infinite. This configuration inhibits any charge transfer between the electrodes, thereby preserving the voltage difference induced by the polarization of the piezoelectric material. Conversely, in the short-circuit condition, the top and bottom electrodes are connected with zero electrical resistance. This allows any charge accumulation due to the polarization of the piezoelectric material to be immediately neutralized, resulting in no electric field within the piezoelectric material. As indicated by the equations (4), the change in the electric field within the piezoelectric material directly affects its stiffness. In practical applications, the device typically operates with a finite electrical resistance connected to electronic circuits, resulting in a state between these two conditions.
Under open-circuit conditions (where $\hat{\bm{q}}=0$), the eigenvalue problem is expressed as follows:
$\begin{split}\bm{C}^{E}\bm{s}(\bm{u}_{oci})-\omega_{oci}^{2}\rho\bm{u}_{oci}+\bm{e}^{T}\bm{\nabla}\varphi_{oci}&=0\\ \bm{e}\bm{s}(\bm{u}_{oci})-\bm{\varepsilon}^{S}\bm{\nabla}\varphi_{oci}&=0.\end{split}$ (5)
Here, $\omega_{oci}$ is the $i$-th mode eigenfrequency under open-circuit conditions, $\bm{u}_{oci}$ is the eigenvector corresponding to the eigenvalue $\omega_{oci}$, and $\varphi_{oci}$ denotes the electric potential for the $i$-th vibration mode under open-circuit conditions. Under open-circuit conditions, the piezoelectric effect increases the stiffness, leading to a higher eigenfrequency $\omega_{oci}$. Conversely, under short-circuit conditions (where $\bm{\nabla}\varphi=0$), the eigenvalue problem reduces to the following equation , unaffected by the influence of the piezoelectric material.
$\bm{C}^{E}\bm{s}(\bm{u}_{sci})-\omega_{sci}^{2}\rho\bm{u}_{sci}=0.$ (6)
Here, $\omega_{sci}$ is the $i$-th mode eigenfrequency under short-circuit conditions, and $\bm{u}_{sci}$ is the eigenvector corresponding to the eigenvalue $\omega_{sci}$. Under short-circuit conditions, the stiffness is reduced due to the neutralization of the piezoelectric effect, resulting in a lower eigenfrequency $\omega_{sci}$.
Comparing the eigenfrequencies under these two conditions provides insight into the impact of the piezoelectric effect on the device’s behavior. The difference in stiffness between open-circuit and short-circuit conditions results in a difference between the eigenfrequencies $\omega_{oci}$ and $\omega_{sci}$. By comparing these eigenfrequencies, the electromechanical coupling coefficient $k^{2}_{i}$ can be calculated, serving as a measure of the effectiveness of the piezoelectric material in converting mechanical energy into electrical energy. The electromechanical coupling coefficient for each vibration mode is given by the following equation [37]:
@@
In the context of a unimorph cantilevered piezoelectric energy harvester design, the eigenfrequency of the device is established as a crucial design parameter for the following reasons. First, these devices generate maximum output when they are resonating at their eigenfrequency, as the energy conversion efficiency is highest in this state. Second, by tuning the eigenfrequency to match the frequency of the ambient vibration source, the harvester can achieve optimal performance at its installation location by effectively capturing and converting the environmental vibration energy into electrical energy [4]. Consequently, in this study, we considered strategies for concurrently aligning the eigenfrequency with the desired value. Adopting the objective function proposed by Yamada et al. [38], we normalized the eigenfrequency specification problem with respect to the $i$-th mode desired frequency $\bar{\omega}_{i}$ and formulated the objective function $F_{\omega}$ as follows:
@@
where $n$ donates the number of modes to evaluation. In this study, for the design of a 3-dimensional beam, we set $n=4$. This is because we need to consider not only the first mode of vibration but also the second mode and bending vibrations in the direction perpendicular to the primary vibration direction, as well as torsional vibrations of the beam, which collectively require evaluating up to the fourth mode.
In this study, we aim to design a high-performance piezoelectric energy harvester by maximizing the electromechanical coupling coefficient $k^{2}_{i}$. To transform the maximization problem into a minimization problem, we formulated the objective function $F_{k}$ using the electromechanical coupling coefficient $k_{i}$ as follows: